OpenStudy (anonymous):

Choose the correct simplification of the expression (2ab^5)^3

OpenStudy (zzr0ck3r):

real numbers are commutative and associative so (2ab5)3 = (2*5*3)(a*b) can you get it from here?

jimthompson5910 (jim_thompson5910):

(2ab^5)^3 = (2^3)*(a^3)*(b^5)^3 What's next?

OpenStudy (anonymous):

add them?

jimthompson5910 (jim_thompson5910):

2^3 = ???

OpenStudy (zzr0ck3r):

did this get edited? it used to say (2ab5)3